OpenStudy (anonymous):

A bank pays 7% interest compounded annually. What principal will grow to $100,000 in 9 years? (Round your answer to two decimal places.)

OpenStudy (anonymous):

$54,393.37

OpenStudy (anonymous):

A = P (1 + r/n)^nt 100,000 = P (1.07)^9 100,000/(1.07)^9 = P

OpenStudy (anonymous):

why cant i take the power of 9 first?